# Transferring Styles for Reduced Texture Bias and Improved Robustness in Semantic Segmentation Networks

### v1

@@ II Related work
Texture bias poses particular challenges in semantic segmentation, where maintaining the precise alignment between input images and segmentation masks is critical. [40] and [41] reported that segmentation models can fail under varying texture conditions even when object shapes are preserved. To better understand the contributions of different visual cues, [42] decomposed images into color, shape, and texture, and [43] confirmed a strong texture bias in models such as SAM [44]. The only other attempt to reduce texture bias directly in semantic segmentation models is training models on images diffused with edge-enhancing diffusion (EED) as used by [17] which leads to a surprsing generalization to the original image domain. Style transfer has been used to analyze properties of instance segmentation models [45] and for debiased semantic segmentation training [46]. Recently, approaches employing neural style transfer for segmentation [47, 48] have shown promising results. In contrast to these methods, for the first time, we measure the effect of style transfer training on shape bias in semantic segmentation employing the evaluation protocol by [16] and evaluate its effect on adversarial attacks directly in semantic segmentation models. In addition, we are the first to apply style transfer at Voronoi image partition level.
@@ IV Experimental setup
For the PASCAL Context dataset, we use the full set of images annotated with semantic classes, splitting it into training and which we partition into validation and test samples to enable further parameter exploration. To preserve shape structure during stylization, we upscale the images to for style transfer and subsequently downscale to for training. All experiments are conducted using the DeepLabV3+ architecture with a ResNet-50 backbone, trained for iterations with a batch size of . We train on original images, on fully AdaIN-stylized images and with Voronoi-based stylization – here we restrict experiments to cells per image due to the lower resolution.
@@ V Numerical results
To investigate whether certain classes suffer or benefit more from stylization, we exemplary compare per-class IoUs of baseline models and the models trained on stylized images which achieved the highest in CDSB for Cityscapes. The result tables with the complete class lists are displayed and discussed in more detail in appendix C. The relative order of class IoUs within model architectures between baseline networks and those trained on stylized data remains almost identical, expect for the regime of the less frequent classes. Furthermore, the rare classes in terms of pixel count also show the highest mIoU performance drop under stylization, except for rider. This behavior is observed for both model architectures. For PASCAL Context, we observe that classes suffer differently from the stylization. While areoplane, boat and sheep lose roughly 20 percent points in IoU, classes like bicycle and diningtable lose less than 3 percent points. It seems that there is no semantic grouping for the classes which suffer less or more by stylization. However, the results for the salient classes generalize across the dataset as observed for train, bus and motorcylce.
@@ Appendix A Tested offline augmentations
Figure 9 displays all training data augmentations used in our Cityscapes experiments. We investigate the parameters and in a hyperparameter study. Note that the number of stylized patches may not exactly correspond to the stylization proportion due to the implementation of partial stylization as a Bernoulli distribution over the Voronoi patches.
@@ Appendix B CDSB component dissection
Table I displays the results for each individual evaluation contributing to the CDSB alongside the achieved mIoU on the Cityscapes validation set. The observed numbers largely confirm the concluded results from the key metric evaluations and highlight that the CDSB measure permits the comparison of networks across performance levels. For example, the networks show a worse performance on Voronoi-shuffled and on data than the -trained networks, but obtain equal or higher shape bias scores nonetheless. Standard deviations are observed to be rather low for most training configurations, which indicates a relatively stable reaction to the applied training data augmentations. As expected, baseline models trained on the original dataset achieve the highest mIoUs on the Voronoi-shuffled dataset, which measures prediction quality achieved only from texture information. In the case of DeepLabV3+ models, it appears that any stylization of training images improves the extrapolation to EED data with the highest gains up to 16 stylized regions. The Voronoi 4 variant appears to exhibit the strongest shape bias but in terms of absolute performance on EED-preprocessed images it is exceeded by the models trained on stylized images with 16 segments. The Voronoi 16 training configuration achieves the highest combined evaluation scores by a small margin over stylized segments, with the Voronoi 4 variant presenting the highest shape bias. Analyzing the results for networks trained on partially stylized images, an increase in mIoU on EED images can be observed, accompanied by a simultaneous decrease in mIoU on Voronoi-shuffled images. This indicates that the measured shape bias and texture bias are complementary properties, as suggested in [16]. In summary, the convolutional network strongly benefits from segment-wise style transferred training images with 16 Voronoi cells showing the best results on Cityscapes. When the images are only partially stylized, the best trade-off between performance and increased shape bias is achieved by stylizing half of the image in the performed experiments.
@@ Appendix C Class-wise analysis
In the following, we investigate on the Cityscapes dataset the performance differences between the baseline model and the models trained on stylized images which achieved the highest in CDSB. Results are listed in Table II. The relative order of class IoUs within model architectures between baseline networks and those trained on stylized data remains mostly identical, with a few exceptions, such as the classes train, bus and motorcycle with lower values. This could be caused by the low frequency of these classes, and bicycles and motorcycles becoming harder to distinguish in the stylized images. A comparison between the model architectures shows that the models have similar class IoUs in most cases. Notably, SegFormer models experience a higher drop in IoU for the classes truck, wall and rider than expected from the drop in mIoU. In contrast, train is significantly more difficult to segment for the convolutional neural network trained with stylization. These differences could stem from the different model encoding mechanisms as well as the randomness of the applied style combined with the low frequency at which the class occurs (cf. the high variance for the class train).
@@
For the PASCAL Context dataset, see Table III, some classes seem to suffer more than other when stylization is applied. As discussed in the main paper, the class sheep looses roughly percent points. This suggests that segmenting a sheep relies strongly on the texture and it is hard to learn their shape features due to texture augmentation. To investigate semantic groups, we split the classes into foreground and background classes to analyze whether foreground objects benefit more or less from stylization. We define following classes as foreground: aeroplane, bicycle, bird, boat, bottle, bus, car cat, chair, cow, diningtable, dog, horse, keyboard, motorbike, person, pottedplant, sheep, sofa, train, tvmonitor. The remaining classes are defined as background. We first observe that the averaged performance of the foreground classes when evaluated on the original data are on par with the averaged performance of the background classes for all trained models (baseline, Voronoi 1, 4, 8 and 16). With respect to the shape bias, we observe an increase in the shape component for an increasing amount of Voronoi cells from 1 to 8. However, the texture component increases simultaneously which in total leads to a drop of the shape bias as observed on mIoU level. For the averaged shape bias over the background classes we observe the contrary behavior for the shape component, i.e., decreases when increasing the amount of differently stylized Voronoi cells. This leads to a generally lower shape bias for the background classes compared to the foreground classes by comparable mIoU performance on the original dataset. Nevertheless, stylization improves the shape biases compared to the baseline for all classes and varying amount of Voronoi cells.
@@ Appendix E Detailed image corruption robustness
The uniform noise distortion adds random values up to the noise intensity to each pixel and color channel. We observe that the SegFormer models are a lot more robust with slower decays in comparison to the DeepLabV3+. The style-transfer-based training image augmentations yield large improvements in robustness for both architectures. The networks trained on partially stylized segments present highly elevated robustness to uniform noise with all tested proportions leading to results similar to those of fully stylized variants. This result corresponds to the observations of increased robustness to adversarial attacks observed in the main manuscript. In the case of SegFormer models, the Voronoi 1 variant shows the largest improvements. Compared to the convolutional network, they appear to be a lot more robust to the introduction of random uniform noise.

### v2

@@ II Related work
Texture bias poses particular challenges in semantic segmentation, where maintaining the precise alignment between input images and segmentation masks is critical. [40] and [41] reported that segmentation models can fail under varying texture conditions even when object shapes are preserved. To better understand the contributions of different visual cues, [42] decomposed images into color, shape, and texture, and [43] confirmed a strong texture bias in models such as SAM [44]. The only other attempt to reduce texture bias directly in semantic segmentation models is training models on images diffused with edge-enhancing diffusion (EED) as used by [17] which leads to a surprising generalization to the original image domain. Style transfer has been used to analyze properties of instance segmentation models [45] and for debiased semantic segmentation training [46]. Recently, approaches employing neural style transfer for segmentation [47, 48] have shown promising results. In contrast to these methods, for the first time, we measure the effect of style transfer training on shape bias in semantic segmentation employing the evaluation protocol by [16] and evaluate its effect on adversarial attacks directly in semantic segmentation models. In addition, we are the first to apply style transfer at Voronoi image partition level.
@@ IV Experimental setup
For the PASCAL Context dataset, we use the full set of images annotated with semantic classes, splitting it into training and , which we partition into validation and test samples to enable further parameter exploration. To preserve shape structure during stylization, we upscale the images to for style transfer and subsequently downscale to for training. All experiments are conducted using the DeepLabV3+ architecture with a ResNet-50 backbone, trained for iterations with a batch size of . We train on original images, on fully AdaIN-stylized images and with Voronoi-based stylization – here we restrict experiments to cells per image due to the lower resolution.
@@ V Numerical results
To investigate whether certain classes suffer or benefit more from stylization, we exemplary compare per-class IoUs of baseline models and the models trained on stylized images which achieved the highest in CDSB for Cityscapes. The result tables with the complete class lists are displayed and discussed in more detail in appendix C. The relative order of class IoUs within model architectures between baseline networks and those trained on stylized data remains almost identical, except for the regime of the less frequent classes. Furthermore, the rare classes in terms of pixel count also show the highest mIoU performance drop under stylization, except for rider. This behavior is observed for both model architectures. For PASCAL Context, we observe that classes suffer differently from the stylization. While areoplane, boat and sheep lose roughly 20 percent points in IoU, classes like bicycle and diningtable lose less than 3 percent points. It seems that there is no semantic grouping for the classes which suffer less or more by stylization. However, the results for the salient classes generalize across the dataset as observed for train, bus and motorcycle.
@@ Appendix A Tested offline augmentations
Figure 9 displays all training data augmentations used in our Cityscapes experiments. We investigate the parameters and in a hyperparameter study. Note that the number of stylized patches may not exactly correspond to the stylization proportion due to the implementation of partial stylization as a Bernoulli distribution over the Voronoi patches. We measured the runtime of our patch-wise style transfer with varying numbers of Voronoi cells on an NVIDIA Quadro P6000 GPU and present the results in Table I. While regular AdaIN is fast, the added Voronoi overhead increases with cell count. This makes the method impractical for online use but suitable for offline precomputation.
| Style Transfer Method | Voronoi Cells | Mean (s) | Std (s) |
| AdaIN (no Voronoi) | 1 | 0.163 | 0.037 |
| Voronoi stylization | 4 | 1.095 | 0.027 |
| 8 | 2.070 | 0.014 | |
| 16 | 4.088 | 0.018 | |
| 32 | 8.168 | 0.019 |
@@ Appendix B CDSB component dissection
Table II displays the results for each individual evaluation contributing to the CDSB alongside the achieved mIoU on the Cityscapes validation set. The observed numbers largely confirm the concluded results from the key metric evaluations and highlight that the CDSB measure permits the comparison of networks across performance levels. For example, the networks show a worse performance on Voronoi-shuffled and on data than the -trained networks, but obtain equal or higher shape bias scores nonetheless. Standard deviations are observed to be rather low for most training configurations, which indicates a relatively stable reaction to the applied training data augmentations. As expected, baseline models trained on the original dataset achieve the highest mIoUs on the Voronoi-shuffled dataset, which measures prediction quality achieved only from texture information. In the case of DeepLabV3+ models, it appears that any stylization of training images improves the extrapolation to EED data with the highest gains up to 16 stylized regions. The Voronoi 4 variant appears to exhibit the strongest shape bias but in terms of absolute performance on EED-preprocessed images it is exceeded by the models trained on stylized images with 16 segments. The Voronoi 16 training configuration achieves the highest combined evaluation scores by a small margin over stylized segments, with the Voronoi 4 variant presenting the highest shape bias. Analyzing the results for networks trained on partially stylized images, an increase in mIoU on EED images can be observed, accompanied by a simultaneous decrease in mIoU on Voronoi-shuffled images. This indicates that the measured shape bias and texture bias are complementary properties, as suggested in [16]. In summary, the convolutional network strongly benefits from segment-wise style transferred training images with 16 Voronoi cells showing the best results on Cityscapes. When the images are only partially stylized, the best trade-off between performance and increased shape bias is achieved by stylizing half of the image in the performed experiments.
@@ Appendix C Class-wise analysis
In the following, we investigate on the Cityscapes dataset the performance differences between the baseline model and the models trained on stylized images which achieved the highest in CDSB. Results are listed in Table III. The relative order of class IoUs within model architectures between baseline networks and those trained on stylized data remains mostly identical, with a few exceptions, such as the classes train, bus and motorcycle with lower values. This could be caused by the low frequency of these classes, and bicycles and motorcycles becoming harder to distinguish in the stylized images. A comparison between the model architectures shows that the models have similar class IoUs in most cases. Notably, SegFormer models experience a higher drop in IoU for the classes truck, wall and rider than expected from the drop in mIoU. In contrast, train is significantly more difficult to segment for the convolutional neural network trained with stylization. These differences could stem from the different model encoding mechanisms as well as the randomness of the applied style combined with the low frequency at which the class occurs (cf. the high variance for the class train).
@@
For the PASCAL Context dataset, see Table IV, some classes seem to suffer more than other when stylization is applied. As discussed in the main paper, the class sheep looses roughly percent points. This suggests that segmenting a sheep relies strongly on the texture and it is hard to learn their shape features due to texture augmentation. To investigate semantic groups, we split the classes into foreground and background classes to analyze whether foreground objects benefit more or less from stylization. We define following classes as foreground: aeroplane, bicycle, bird, boat, bottle, bus, car cat, chair, cow, diningtable, dog, horse, keyboard, motorbike, person, pottedplant, sheep, sofa, train, tvmonitor. The remaining classes are defined as background. We first observe that the averaged performance of the foreground classes when evaluated on the original data are on par with the averaged performance of the background classes for all trained models (baseline, Voronoi 1, 4, 8 and 16). With respect to the shape bias, we observe an increase in the shape component for an increasing amount of Voronoi cells from 1 to 8. However, the texture component increases simultaneously which in total leads to a drop of the shape bias as observed on mIoU level. For the averaged shape bias over the background classes we observe the contrary behavior for the shape component, i.e., decreases when increasing the amount of differently stylized Voronoi cells. This leads to a generally lower shape bias for the background classes compared to the foreground classes by comparable mIoU performance on the original dataset. Nevertheless, stylization improves the shape biases compared to the baseline for all classes and varying amount of Voronoi cells.
@@ Appendix E Detailed image corruption robustness
Uniform noise distortion adds random values up to the noise intensity to each pixel and color channel. We observe that the SegFormer models are a lot more robust with slower decays in comparison to the DeepLabV3+. The style-transfer-based training augmentations yield large improvements in robustness for both architectures. Networks trained on partially stylized segments present highly elevated robustness to uniform noise with all tested proportions leading to results similar to those of fully stylized variants. This corresponds to the observations of increased robustness to adversarial attacks observed in the main manuscript. In the case of SegFormer models, the Voronoi 1 variant shows the largest improvements. Compared to the CNN, they are a lot more robust to the introduction of random uniform noise.